3. On January 1, 2017, Chic Corp paid $750,000 for 100.000 shares of Toto Companted common stock, which represents 25% of Totos outstanding common st ock. Toto reported f$300. 000 and paid cash dividends of $80.000 during 2017 Chic should report in Toto Company on its December 31, 2017, balance sheet at a. $750,000 b. $825,000 c $770,000 d. $805,000 57 CA 4. McComb Inc. earns $1,350,000 and pays cash dividends for $450,000 during 2017. SFX Corporation owns 70,000 of the 210,000 outstanding shares of McComb. What amount should SFX show in the investment account at December 31, 2017 if the beginning of the year balance in the account was $150,000? a. $450,000 b. $300,000 C. $420,000 d. $600,000 59 McComb Inc. earns $1,350,000 and pays cash dividends for $450,000 during 2017. SFX Corporation owns 70,000 of the 210,000 outstanding shares of McComb. How much revenue from investment should SFX report in 2017? a. $150,000 b. $300,000 C. $450,000 5. d. $600,000 All of the following factors would be signs of an investors significant influence over an investee except a. the investor has representation on the investees board of directors. b. the investor participates in the investees policy-making process. c. there are immaterial transactions between the investor and the investee. d. the common stock held by other stockholders is dispersed. APPENDIX 1. . Which one of the following payroll taxes is not withheld from the employees wages because it is not levied on the employee? a. Federal income tax b. Federal unemployment tax c. State income tax d. FICA tax 2.. The tax that is paid equally by the employer and employee is the a. federal income tax. b. federal unemployment tax. c. state unemployment tax d. FICA tax.
Answer #1
3 Option D is the correct answer
The Equity Method
Investment in Toto Company
Purchase 750000
Retained Earnings 75000 =300000*25%
Less: Dividend -20000 =-80000*25%
805000
4 Option A is the correct answer
Investment in McComb Inc
Investment in McComb Inc 33.33% =70000/210000
The Equity Method
Investment in McComb Inc
Beginning 150000 150000
Retained Earnings 450000 =1350000*(70000/210000)
Less: Dividend -150000 =-450000*(70000/210000)
450000
5 Option C is the correct answer
Investment in McComb Inc 33.33% =70000/210000
Revenue from Investment 450000 =1350000*(70000/210000)
6 Option C is the correct answer